An orbiting satellite has seven solar cells, which function independently of each other. During a mission, each cell fails with a probability of 0.3. What is the probability that there will be adequate power output for the mission if 7. All 7 cells must be active?

Verified Answer :

toporc

The probability of each cell not failing is 1.0 - 0.3 = 0.7.
The probability of all 7 cells not failing during a mission is given by:
\(P(7\ cells\ do\ not\ fail)=0.7^{7}=0.082\ or\ 8.2\ percent\)